A certain reaction is spontaneous at 90.°C. If the enthalpy change for the reaction is 30. kJ, what is the minimum value of

Δr S (in joules per kelvin) for the reaction?


1
Expert's answer
2020-12-10T07:49:24-0500

ΔS=-ΔH/T

T= (90+273) =363K

ΔS=-30000/363

  =-82.64J/K